# Engineering a Josephson junction chain for the simulation of the clock model

### v1

@@ Abstract
The continuous improvement of fabrication techniques and high-quality semiconductor-superconductor interfaces allowed for unprecedented tunability of Josephson junction arrays (JJA), making them a promising candidate for analog quantum simulations of many-body phenomena. While most experimental proposals so far focused on quantum simulations of ensembles of two-level systems, the possibility of tuning the current-phase relation beyond the sinusoidal regime paves the way for studying statistical physics models with larger local Hilbert spaces. Here, we investigate a particular JJA architecture that can be mapped into a $\mathbb{Z}_{3}$ clock model. Through matrix-product-states simulations and bosonization analysis, we show that few experimentally accessible control parameters allow for the exploration of the rich phase diagrams of the associated low-energy field theories. Our results expand the horizon for analog quantum simulations with JJAs towards models that can not be efficiently captured with qubit architectures.
@@ I Introduction
Here we propose a solid-state realization of quantum clock models in one dimension based on hybrid semiconductor / superconductor devices [28, 29]. These platforms combine the coherence of superconducting systems with the tunability via electrostatic gates offered by their semiconducting substrate [29, 30, 31, 32]. Moreover, their fabrication techniques allow for a considerable scalability, with systems of thousands of Josephson junctions studied in recent experiments [33, 34, 35].
These experimental architectures offer an ideal platform to study in a controllable way fundamental quantum many-body phenomena, as BKT transitions [33, 36, 37], and to investigate the properties of emergent quantum field theories [38, 39], including integrable [40] and conformal models [41, 42].
In this work, we take advantage of their tunability to design superconducting chains with an emergent $\mathbb{Z}_{3}$ symmetry. The platform that we design presents a very rich phase diagram that enables the observation of non-trivial quantum critical states including conformal Potts phase transitions, tricritical Lifshitz lines, and gapless floating phases [14, 15]. A few experimentally accessible global parameters span this phase diagram, which we characterize through a combination of field-theory approaches based on bosonization and matrix-product-states (MPS) numerical techniques optimized with the density matrix renormalization group (DMRG) algorithm.
Our construction is based on a building block that relies on the non-sinusoidal energy-phase relation typical of hybrid Josephson junctions. In Sec. II we describe a triple Josephson junction that can be operated as a quantum clock degree of freedom when tuned in a trasmon-like regime. Based on these elementary units, we derive the Hamiltonian for the Josephson junction array (JJA) in Sec. III and its mapping on $\mathbb{Z}_{3}$ clock models, while in Sec. IV we discuss its low-energy effective field-theory description. Section V contains the bulk of our numerical results. We first analyze the Potts phase transition in Sec. V.1 and its classical limit in Sec. V.2, then we characterize the phase diagram of the chiral clock model in Sec. V.3. In Sec. VI we describe the role played by the induced charge on the SC islands and we discuss the deviations from the $\mathbb{Z}_{3}$-symmetric case and the effect of disorder in Sec. VII. Finally, we draw our conclusions and discuss possible outlooks in Sec. VIII. Further analytical and numerical analysis are presented in the Appendix.
@@ II A triple Josephson junction operated as a quantum clock
Figure 1: Sketch of the JJA used to simulate the $\mathbb{Z}_{3}$ clock Hamiltonian. Each SC island (blue) is coupled through tunable Josephson junctions to its two neighbors and a common grounded SC lead.
The first task to design a one-dimensional Josephson junction chain for the quantum simulation of models with a global $\mathbb{Z}_{3}$ symmetry is the definition of a suitable elementary block that can be operated as a 3-state quantum clock (or, equivalently, as a qutrit). Therefore we must consider an elementary superconducting (SC) element that, in the ideal limit, displays three degenerate ground states, which we label by $|s\rangle$ with $s=0,1,2$. It is useful to introduce a pair of clock operators, $\sigma$ and $\tau$, acting on these three states in the following way:
@@
where we are assuming that the Cooper pair tunneling is mediated by a single Andreev state; $10<T_{p}<$ 1 represents the transparency of the junction and $\Delta$ is the SC gap induced in the semiconducting 2DEG. Recent studies have shown that a similar dispersion can also be achieved by considering two traditional junctions in series [44, 45]. Importantly, the parameter $T_{p}$ can be tuned via suitable electrostatic gates in hybrid semiconductor / superconductor devices [30, 31, 32] and, in turn, it determines the ratio between the harmonics $\cos n\varphi$ appearing in $\mathcal{E}_{J}$.
To obtain the ideal potential (3), we consider a device with three identical Josephson junctions, such that $T_{1}=T_{2}=T_{3}$, arranged to form two SC loops with the same area (Fig. 1). Through an out-of-plane magnetic field, we introduce the same magnetic flux $\Phi$ in both loops, such that $\Phi/\Phi_{0}=1/3$, where $\Phi_{0}=hc/2e$ is the magnetic flux quantum. Under these conditions, analogously to Ref. [42], we obtain that the total potential $\mathcal{E}_{J}^{(1)}\left(\varphi-\frac{2\pi}{3}\right)+\mathcal{E}_{J}^{(2)}\left(\varphi\right)+\mathcal{E}_{J}^{(3)}\left(\varphi+\frac{2\pi}{3}\right)$ assumes the form in Eq. (3). The requirement $T_{1}=T_{2}=T_{3}$ can be achieved in experiments by tuning the external transparencies $T_{1}$ and $T_{3}$ via suitable electrostatic gates. Perturbations around this optimal point will result in the splitting of the minima of the potential, which we will address in Sec. VII. A similar construction was considered in Ref. [46], where a so-called masus building block was designed to obtain the potential (3) with three SC arms in parallel, each composed by two sinusoidal Josephson junctions.
@@
To understand the physics of the triple Josephson junction we can consider different limits. When ${J_{0}}$ dominates over $E_{c}$, this system lies in a transmon-like regime: the three degenerate minima of the potential $V(\varphi)$ determine the three ground states $|s\rangle$ whose wavefunctions are peaked around $\varphi=0,2\pi/3$ and $34\pi/$ 3. Hence, the triple junction represents a quantum clock degree of freedom labelled by the operator $\sigma\sim\mathrm{e}^{i\varphi}$. The charging energy plays the role of a kinetic term which causes phase slips between these three ground states. Analogously to transmons [47], the related amplitudes can be estimated based on a semiclassical approximation and the effective low-energy Hamiltonian corresponds to the following description in terms of the $\tau$ clock operator:
$H_{\rm eff}\overset{{J_{0}}>E_{c}}{\approx}-\alpha E_{c}^{1/4}{J_{0}}^{3/4}\mathrm{e}^{-\sqrt{\frac{32{J_{0}}}{9E_{c}}}+in_{g}\frac{2\pi}{3}}\tau+{\rm H.c.}\\ \equiv-h_{\tau}\left(e^{in_{g}\frac{2\pi}{3}}\tau+\mathrm{e}^{-in_{g}\frac{2\pi}{3}}\tau^{\dagger}\right)\,,$ (6)
@@
For increasing $E_{c}$, however, the splitting of the three lowest-energy states caused by the phase slips becomes of order $\sim E_{c}$ and the three-level approximation fails if $E_{c}\sim{J_{0}}$. By increasing $E_{c}$ further, the triple junction enters a Coulomb-dominated regime ($E_{c}\gg{J_{0}}$); its eigenstates evolve toward charge eigenstates with energies approximately described by the charging energy term. The triple-junction maintains its $\mathbb{Z}_{3}$ symmetry, since the total number of Cooper pairs is conserved modulo 3. However it cannot be described any longer as a clock degree of freedom.
@@ III The Josephson junction chain
To design a system with a global $\mathbb{Z}_{3}$ symmetry, we consider a chain of SC islands connected to the same background superconductor via the triple Josephson junctions described in the previous section (see Fig. 1), which provides a specific example of the models presented in Ref. [41]. Each SC island is coupled with its neighbors through a standard sinusuoidal Josephson junction of energy $E_{J}$ and the related Hamiltonian reads:
@@
Finally, we observe that in uniform systems with position-independent phases $A$, both Hamiltonians (7) and (9) fulfill the following properties:
@@ IV Effective field theory description of the model
Additional electrostatic interactions among neighboring islands would influence the parameters $K,v$ and $M$ without qualitatively affecting the field Hamiltonian (10).
@@ V Observables and phase diagram
In this truncated space, the creation and annihilation operators of the Cooper pairs, $\mathrm{e}^{-i\varphi_{j}}$ and $\mathrm{e}^{i\varphi_{j}}$, are respectively approximated by off-diagonal $(2N_{\rm max}+1)\times(2N_{\rm max}+1)$ matrices $\Sigma_{j}^{+}$ and $\Sigma_{j}^{-}$, which raise and lower the particle number of the island $j$ (see, for instance, Ref. [42]). We also observe that, in the large ${J_{0}}$ limit, where the mapping to the clock model (5) is most accurate, the matrices $\Sigma_{j}^{+}$ and $\Sigma_{j}^{-}$ respectively converge to the clock operators $\sigma^{\dagger}_{j}$ and $\sigma_{j}$. Therefore the correspondence between phase operators, truncated matrices for their description and clock operators is given by:
@@ V.1 The Potts model regime: $n_{g}=0$, $A=0$
Figure 2: Summary of results obtained with DMRG: ground state energy, long-range order parameter, total charge fluctuations, correlation length. Comparison between four system sizes $L$ with a charge cutoff of $N_{\rm max}=7$, corresponding to 15 different charge states for each SC island. The other physical parameters are $E_{J}=0.2{J_{0}}$, $n_{g}=0$, and $A=0$. The vertical in panels (b) and (c) mark the phase transition occurring at $E_{c}=E_{c}^{*}\simeq 1.62E_{J}$. The dashed black lines are power-law fit giving critical exponents $\beta=0.089(2)$, $\nu^{+}=0.88(5)$ for $E_{c}>E_{c}^{*}$ and $\nu^{-}=0.87(6)$ for $E_{c}<E_{c}^{*}$.
@@ V.2 The chiral classical limit: $E_{c}\to 0$, $A\neq 0$
Given the symmetries of the system, we restrict our analysis to $30\leq A\leq 2\pi/$ 3. The classical configurations with minimal energy are depicted in Fig. 4 and the boundaries between the commensurate and the incommensurate classical ground states is easily found by comparing their average energy densities: the energy density of the ferromagnetic configuration is given by $\varepsilon=-{J_{0}}-E_{J}\cos A$; the one for the incommensurate phase with winding $A$ is instead $-E_{J}$, since the onsite potential averages to zero. Their boundary is thus given by:
@@
Figure 4: Phase diagram of the classical limit ($E_{c}=0$) of the Josephson junction chain. Depending on $A$ and the ratio $J_{0}/E_{J}$ the classical ground state corresponds to the ferromagnetic phase ($w=0$), the helical phase ($w=1$), or the incommensurate phase ($\varphi_{j}=Aj$).
@@
This exponential degeneracy is split when leaving the classical limit by introducing a weak charging energy $E_{c}>0$, thus $h_{\tau}>0$. The behavior of the system for $A\sim\pi/3$ and ${J_{0}}>E_{J}/2$ can be described based on a mapping of the chiral clock into a system of free fermions presented by Ostlund [14]. The link variable $w_{j}=0,1$ can indeed be considered as the occupation number of a hardcore boson since link configurations of the clock model with $w_{j}=2$ are suppressed by an energy gap $23E_{J}/$ 2, playing the role of an onsite interaction term of the bosonic degrees of freedom. These hardcore bosons can be mapped into fermionic degrees of freedom by a Jordan-Wigner transformation. Additionally, the transverse field term $h_{\tau}$ in Eq. (9) becomes a kinetic energy term for these fermions. The model for $A=\pi/3$, ${J_{0}}>E_{J}/2$ and $E_{c}\ll E_{J}$ can thus be mapped into a system of fermions at half filling with bandwidth $2h_{\tau}$.
@@
This shift plays the role of a chemical potential for the hardcore boson description. Finally, for weak $h_{\tau}\ll E_{J}$, at second order in perturbation theory, a further term emerges that can annihilate three neighboring bosons $w_{j}=1$ by rotating two rotors. Overall, the model at low $E_{c}$ in proximity with $A=\pi/3$ can thus be described by:

### v2

# Engineering a Josephson junction chain for the simulation of the quantum clock model

@@ Abstract
The continuous improvement of fabrication techniques and high-quality semiconductor-superconductor interfaces allows for an unprecedented tunability of Josephson junction arrays (JJA), making them a promising candidate for analog quantum simulations of many-body phenomena. While most experimental proposals so far focused on quantum simulations of ensembles of two-level systems, the possibility of tuning the current-phase relation beyond the sinusoidal regime paves the way for studying statistical physics models with larger local Hilbert spaces. Here, we investigate a particular JJA architecture that can be mapped into a $\mathbb{Z}_{3}$ clock model. Through matrix-product-states simulations and bosonization analysis, we show that few experimentally accessible control parameters allow for the exploration of the rich phase diagrams of the associated low-energy field theories. Our results expand the horizon for analog quantum simulations with JJAs towards models that can not be efficiently captured with qubit architectures.
@@ I Introduction
Here we propose a solid-state realization of quantum clock models in one dimension based on hybrid semiconductor-superconductor devices [28, 29]. These platforms combine the coherence of superconducting (SC) systems with the tunability via electrostatic gates offered by their semiconducting substrate [29, 30, 31, 32]. Moreover, their fabrication techniques allow for considerable scalability, with systems of thousands of Josephson junctions studied in recent experiments [33, 34, 35].
These experimental architectures offer an ideal platform to explore in a controllable way fundamental quantum many-body phenomena, as BKT transitions [33, 36, 37], and to investigate the properties of emergent quantum field theories [38, 39], including integrable [40] and conformal models [41, 42].
In this work, we take advantage of their tunability to design SC chains with an emergent $\mathbb{Z}_{3}$ symmetry. The platform that we design presents a very rich phase diagram that enables the observation of non-trivial quantum critical states including conformal Potts phase transitions, tricritical Lifshitz lines, and gapless floating phases [14, 15]. A few experimentally accessible global parameters span this phase diagram, which we characterize through a combination of field-theory approaches based on bosonization and matrix-product-states (MPS) numerical techniques optimized with the density matrix renormalization group (DMRG) algorithm.
Our construction is based on a building block that relies on the non-sinusoidal energy-phase relation typical of hybrid Josephson junctions. In Sec. II we describe a triple Josephson junction that can be operated as a quantum clock degree of freedom when tuned in a transmon-like regime. Based on these elementary units, we derive the Hamiltonian for the Josephson junction array (JJA) in Sec. III and its mapping on $\mathbb{Z}_{3}$ clock models, while in Sec. IV we discuss its low-energy effective field-theory description. Section V contains the bulk of our numerical results. We first analyze the Potts phase transition in Sec. V.1 and its classical limit in Sec. V.2, then we characterize the phase diagram of the chiral clock model in Sec. V.3. In Sec. VI we describe the role played by the induced charge on the SC islands and we discuss the deviations from the $\mathbb{Z}_{3}$-symmetric case and the effect of disorder in Sec. VII. Finally, we draw our conclusions and discuss possible outlooks in Sec. VIII. Further analytical and numerical analysis are presented in the Appendix.
@@ II A triple Josephson junction operated as a quantum clock
Figure 1: Sketch of the JJA used to simulate the $\mathbb{Z}_{3}$ clock Hamiltonian. Each SC island (blue) is coupled to a common grounded SC background through three parallel tunable Josephson junctions with transparencies $T_{p}$. Cooper pairs can additionally tunnel between neighboring islands via sinusoidal Josephson junctions with energy $E_{J}$.
The first task to design a one-dimensional Josephson junction chain for the quantum simulation of models with a global $\mathbb{Z}_{3}$ symmetry is the definition of a suitable elementary block that can be operated as a 3-state quantum clock (or, equivalently, as a qutrit). Therefore we must consider an elementary SC element that, in the ideal limit, displays three degenerate ground states, which we label by $|s\rangle$ with $s=0,1,2$. It is useful to introduce a pair of clock operators, $\sigma$ and $\tau$, acting on these three states in the following way:
@@
where we are assuming that the Cooper pair tunneling is mediated by a single Andreev state; $10<T_{p}<$ 1 represents the transparency of the junction and $\Delta$ is the SC gap induced in the semiconducting 2DEG.
To engineer the ideal potential (3), we consider a device with three identical Josephson junctions, such that $T_{1}=T_{2}=T_{3}$, arranged to form two SC loops with the same area (Fig. 1).
Through an out-of-plane magnetic field, we introduce the same magnetic flux $\Phi$ in both loops, such that $\Phi/\Phi_{0}=1/3$, where $\Phi_{0}=hc/2e$ is the magnetic flux quantum. Under these conditions, analogously to Ref. [42], we obtain that the total potential $\mathcal{E}_{J}^{(1)}\left(\varphi-\frac{2\pi}{3}\right)+\mathcal{E}_{J}^{(2)}\left(\varphi\right)+\mathcal{E}_{J}^{(3)}\left(\varphi+\frac{2\pi}{3}\right)$ can be approximated by the form in Eq. (3). Perturbations around this optimal point will result in the splitting of the minima of the potential, which we will address in Sec. VII.
Concerning the experimental realization of this symmetric triple junction, we emphasize that the general scenario with Josephson junctions characterized by multiple Andreev states does not yield qualitative differences to implement the potential (3): the addition of fluxes $\Phi_{0}/3$ makes indeed the potential periodic in $\varphi$ with periodicity $32\pi/$ 3, with a dominant $\cos 3\varphi$ component. In this respect, more general descriptions of the energy-phase relation (4) can be adopted [43, 44] without affecting the required symmetries.
Josephson junctions with the dispersion $\mathcal{E}_{J}^{(p)}$ can also be realized with alternative techniques. For instance, recent studies have shown that the energy-phase relation (4) can be efficiently achieved by considering two traditional junctions in series [45, 46]. A key aspect is that in hybrid semiconductor-superconductor platforms, their transparency $T_{p}$ can be tuned via suitable electrostatic gates [30, 31, 32], and, in turn, it determines the ratio between the harmonics $\cos n\varphi$ appearing in $\mathcal{E}_{J}$.
The resulting construction of 3-state quantum clocks based on three parallel SC arms, each composed of two sinusoidal Josephson junctions, is reminiscent of the so-called masus devices [47]. The controllability of the external transparencies $T_{1}$ and $T_{3}$ makes it possible to implement the potential (3), and, very recently, their two-arm counterparts, commonly called Josephson rhombi [48], have been successfully realized in tunable hybrid InAs - Al devices [49].
@@
To understand the physics of the triple Josephson junction we can consider different limits. When ${J_{0}}$ dominates over $E_{c}$, this system lies in a transmon-like regime: the three degenerate minima of the potential $V(\varphi)$ determine the three ground states $|s\rangle$ whose wavefunctions are peaked around $\varphi=0,2\pi/3$ and $34\pi/$ 3. Hence, the triple junction represents a quantum clock degree of freedom labeled by the operator $\sigma\sim\mathrm{e}^{i\varphi}$. The charging energy plays the role of a kinetic term which causes phase slips between these three ground states. Analogously to transmons [50], the related amplitudes can be estimated based on a semiclassical approximation and the effective low-energy Hamiltonian corresponds to the following description in terms of the $\tau$ clock operator:
$H_{\rm eff}\overset{{J_{0}}>E_{c}}{\approx}-\alpha E_{c}^{1/4}{J_{0}}^{3/4}\mathrm{e}^{-\sqrt{\frac{32{J_{0}}}{9E_{c}}}+in_{g}\frac{2\pi}{3}}\tau+{\rm H.c.}\\ \equiv-h_{\tau}\left(\mathrm{e}^{in_{g}\frac{2\pi}{3}}\tau+\mathrm{e}^{-in_{g}\frac{2\pi}{3}}\tau^{\dagger}\right)\,,$ (6)
@@
For increasing $E_{c}$, however, the splitting of the three lowest-energy states caused by the phase slips becomes of order $\sim E_{c}$ and the three-level approximation fails if $E_{c}\sim{J_{0}}$. By increasing $E_{c}$ further, the triple junction enters a Coulomb-dominated regime ($E_{c}\gg{J_{0}}$); its eigenstates evolve toward charge eigenstates with energies approximately described by the charging energy term. The triple-junction maintains its $\mathbb{Z}_{3}$ symmetry since the total number of Cooper pairs is conserved modulo 3. However, it cannot be described any longer as a clock degree of freedom.
@@ III The Josephson junction chain
To design a system with a global $\mathbb{Z}_{3}$ symmetry, we consider a chain of SC islands connected to the same background superconductor via the triple Josephson junctions described in the previous section (see Fig. 1), which provides a specific example of the models presented in Ref. [41]. Each SC island is coupled with its neighbors through a standard sinusoidal Josephson junction of energy $E_{J}$ and the related Hamiltonian reads:
@@
The introduction of Josephson junctions between neighboring islands also favors the onset of mutual electrostatic interactions between the islands of the form $4e^{2}(C^{-1})_{j,j^{\prime}}\left(N_{j}-n_{g}\right)\left(N_{j^{\prime}}-n_{g}\right)$, where $(C^{-1})_{j,j^{\prime}}$ are the off-diagonal elements of the inverse capacitance matrix of the system [51]. In Eq. (7) these interactions are neglected: we expect indeed that the self-capacitance between each island and the grounded SC background is considerably stronger than the mutual capacitance between the islands. In particular, Ref. [52] investigated a setup with two transmons connected to the same SC background and linked by a tunable hybrid Josephson junction. In that analysis the maximal ratio between the diagonal terms $(C^{-1})_{j,j}$ and the off-diagonal nearest-neighbor term $(C^{-1})_{j,j+1}$ was about $1/9$ and it was strongly suppressed when increasing the filling the semiconducting region between the two neighboring superconducting islands as effect of the screening provided by the two-dimensional electron gas. The role of inter-island interactions, however, does not qualitatively affect the main features of the system and we will discuss it in Sec. VII.
@@
Finally, we observe that in uniform JJAs with position-independent phases $A$, both Hamiltonians (7) and (9) fulfill the following properties:
@@ IV Effective field theory description of the model
In systems with sizeable mutual interactions between neighboring islands, these electrostatic repulsions further enhance the role of $E_{c}$, thus lowering $K$ and increasing $v$.
@@
Additional electrostatic interactions among neighboring islands would influence the parameters $K,v$, and $M$ without qualitatively affecting the field Hamiltonian (10).
@@ V Observables and phase diagram
In this truncated space, the creation and annihilation operators of the Cooper pairs, $\mathrm{e}^{-i\varphi_{j}}$ and $\mathrm{e}^{i\varphi_{j}}$, are respectively approximated by off-diagonal $(2N_{\rm max}+1)\times(2N_{\rm max}+1)$ matrices $\Sigma_{j}^{+}$ and $\Sigma_{j}^{-}$, which raise and lower the particle number of the island $j$ (see, for instance, Ref. [42]). We also observe that, in the large ${J_{0}}$ limit, where the mapping to the clock model (5) is most accurate, the matrices $\Sigma_{j}^{+}$ and $\Sigma_{j}^{-}$ respectively converge to the clock operators $\sigma^{\dagger}_{j}$ and $\sigma_{j}$. Therefore the correspondence between phase operators, truncated matrices for their description, and clock operators is given by:
@@ V.1 The Potts model regime: $n_{g}=0$, $A=0$
Figure 2: Summary of results obtained with DMRG: ground state energy, long-range order parameter, total charge fluctuations, correlation length. The data display the comparison between four system sizes $L$ with a charge cutoff of $N_{\rm max}=7$, corresponding to 15 different charge states for each SC island. The other physical parameters are $E_{J}=0.2{J_{0}}$, $n_{g}=0$, and $A=0$. The vertical dashed lines in panels (b) and (d) mark the phase transition occurring at $E_{c}=E_{c}^{*}\simeq 1.62E_{J}$. The dashed black lines are power-law fits returning critical exponents $\beta=0.089(2)$, $\nu^{+}=0.88(5)$ for $E_{c}>E_{c}^{*}$ and $\nu^{-}=0.87(6)$ for $E_{c}<E_{c}^{*}$.
@@ V.2 The chiral classical limit: $E_{c}\to 0$, $A\neq 0$
Given the symmetries of the system, we restrict our analysis to $30\leq A\leq 2\pi/$ 3. The classical configurations with minimal energy are depicted in Fig. 4 and the boundaries between the commensurate and the incommensurate classical ground states are easily found by comparing their average energy densities: the energy density of the ferromagnetic configuration is given by $\varepsilon=-{J_{0}}-E_{J}\cos A$; the one for the incommensurate phase with winding $A$ is instead $-E_{J}$, since the onsite potential averages to zero. Their boundary is thus given by:
@@
Figure 4: Phase diagram of the classical limit ($E_{c}=0$) of the Josephson junction chain. Depending on $A$ and the ratio $J_{0}/E_{J}$ the classical ground state corresponds to the ferromagnetic phase ($w=0$), the helical phase ($w=1$), or the incommensurate (IC) phase ($\varphi_{j}=Aj$).
@@
This exponential degeneracy is split when leaving the classical limit by introducing a weak charging energy $E_{c}>0$, thus $h_{\tau}>0$. The behavior of the JJA for $A\sim\pi/3$ and ${J_{0}}>E_{J}/2$ can be described based on a mapping of the chiral clock into a system of free fermions presented by Ostlund [14]. The link variable $w_{j}=0,1$ can indeed be considered as the occupation number of a hardcore boson since link configurations of the clock model with $w_{j}=2$ are suppressed by an energy gap $23E_{J}/$ 2, playing the role of an onsite interaction term of the bosonic degrees of freedom. These hardcore bosons can be mapped into fermionic degrees of freedom by a Jordan-Wigner transformation. Additionally, the transverse field term $h_{\tau}$ in Eq. (9) becomes a kinetic energy term for these fermions. The model for $A=\pi/3$, ${J_{0}}>E_{J}/2$ and $E_{c}\ll E_{J}$ can thus be mapped into a system of fermions at half filling with bandwidth $2h_{\tau}$.
@@
This shift plays the role of a chemical potential for the hardcore boson description. Finally, for weak $h_{\tau}\ll E_{J}$, at second order in perturbation theory, a further term emerges that can annihilate three neighboring bosons $w_{j}=1$ by rotating two rotors. Overall, the model at low $E_{c}$ in proximity of $A=\pi/3$ can thus be described by: